OpenStudy (anonymous):

Jerry typed 185 words in 6 minutes. Pam typed 125 words in 4 minutes. Sheryl typed 242 words in 8 minutes. Tom typed 354 words in 12 minutes. Who typed at the fastest rate?

OpenStudy (zzr0ck3r):

divide them

OpenStudy (zzr0ck3r):

185/6 will show how much the first person is typing per minute.

OpenStudy (anonymous):

Tom for sure

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

wrong

OpenStudy (anonymous):

:o

OpenStudy (anonymous):

Tom was actually the slowest lol

OpenStudy (anonymous):

Type it out for me like @zzr0ck3r said

OpenStudy (anonymous):

oh ur suppost to choose the like least? cuz somebody had 500 but I dont remeber who maybe Pam?

OpenStudy (anonymous):

185/6 = ? 125/4 = ? 242/8 = ? 354/12 = ?

OpenStudy (zzr0ck3r):

\(\uparrow \) pick the biggest one

OpenStudy (anonymous):

are you finding the fastest or slowest typed ? @Lindsay_552

OpenStudy (anonymous):

fastest

OpenStudy (anonymous):

okay so which one has the biggest answer

OpenStudy (anonymous):

witch would be Tom right

OpenStudy (anonymous):

cuz he has 4,140

OpenStudy (anonymous):

No what are doing !

OpenStudy (anonymous):

divide each one and tell me what you get

OpenStudy (anonymous):

OH I WAS MULTIPLYING

OpenStudy (anonymous):

OMG !

OpenStudy (anonymous):

you still didnt do what I asked you are dividing

OpenStudy (anonymous):

Answer this now : 185/6 = ? 125/4 = ? 242/8 = ? 354/12 = ?

OpenStudy (zzr0ck3r):

divide the numbers...... \(\frac{185}{6}\approx 30.8333\)

OpenStudy (anonymous):

185/6= 30.83333333 125/4= 31.25 242/8= 30.25 354/12= 29.5

OpenStudy (anonymous):

Ok so now what's the answer

OpenStudy (anonymous):

who is the fastest

OpenStudy (anonymous):

Pam

OpenStudy (anonymous):

Yes finally !

OpenStudy (anonymous):

lol told u im stupid

OpenStudy (anonymous):

my twin is the smart one

OpenStudy (anonymous):

lol no you just don't listen

OpenStudy (anonymous):

my parents say that all the time im to stubbern to listen

OpenStudy (anonymous):

they say

OpenStudy (anonymous):

They must be right ^_^

OpenStudy (anonymous):

lol Lynden listens to everything probs why he is so darn smart u know I try to have him help me he NEVER does so I come on here

OpenStudy (anonymous):

Lol wow

OpenStudy (anonymous):

He should be helping you

OpenStudy (anonymous):

I have 3 brothers I'm the only girl its so sad :( lol now im being random

OpenStudy (anonymous):

Lol random is good i'm random

OpenStudy (anonymous):

But next question because unlike you I don't have all night lol I have work

OpenStudy (anonymous):

lol I'm really weird man but u wont believe it my mom is pregnant :o isnt 5 enough!

OpenStudy (anonymous):

guess not and ok one more question

OpenStudy (anonymous):

Lol what's she having a boy ?

OpenStudy (anonymous):

she actually finds out in a few days at her doctors appointment god my figures are crossed for a sister I have two older brothers and then my twin

OpenStudy (anonymous):

Lol yea hopefully she has a girl

OpenStudy (anonymous):

I beg they stop at this baby I mean 5 kids is so much lol

OpenStudy (anonymous):

anyway next question ill post it hang on